# research-article-2col

### v1

@@ 2 Material and Methods
[…]
@@
[…]
Figure 1: Pulse sequence diagram for 3D New AMRF. The overall pulse scheme is shown in (A). The diagram in (B) represents the variable flip angle (FA) and phase angle (PA) MRF train. There were a total of NN = 23 non-selective rectangular RF pulses of duration $\tau_{RF}=0.8$ ms. Each pulse was followed by a time period $\tau_{i}$. On the RF channel in (A), the gray rectangles correspond to a single FA = $\theta_{i}$ and PA = $\phi_{i}$ combination in train (B). In (B), the dark gray rectangles represent a 3-pulse inversion composite block (25) used to increase $\text{T}_{\text{1}}$ sensitivity of the sequence and improve RF homogeneity for the magnetization inversion. Within the composite block, $\tau_{1}$ = $\tau_{2}$ = 7.5 ms. The light gray rectangles indicate the 20-pulse variable FA and PA train. All FA = $\theta_{i}$ and PA = $\phi_{i}$ are listed at the top of each rectangle and $\tau_{i}$ = 15 ms for $i\geq 3$. During each delay $\tau_{i}$, a time period TE = 0.2 ms was followed by the ADC event, indicated by a yellow block on the ADC channel. A 3D spiral encoding scheme using the FLORET trajectory (20, 24, 26) (3 hubs at 45, 100 interleaves/hub), indicated in blue, was played over the gradient channels. Immediately after the ADC, a rewind gradient was also played out and then the next RF pulse in the MRF train was initiated. After $N$ pulses and $N\tau_{i}$ delays have played out, a 5-ms spoiling gradient at 70% maximum gradient strength, indicated by an orange trapezoid, was applied in all directions. The entirety of this scheme represented one TR of the sequence.
@@
Figure 1(A) presents the New AMRF sequence. System excitation was driven by a series of $N$ non-selective rectangular RF pulses with FA $\theta_{i}$ and PA $\phi_{i}$ followed by a delay $\tau_{i}$ (i=1,2,,Ni=1,2,...,N), forming the MRF pulse train shown in Figure 1(B). An initial magnetization inversion using a 90–180–90 (dark gray) composite 3-pulse block (25) was used to increase $\text{T}_{\text{1}}$ sensitivity and improve RF homogeneity for the inversion before initiating the 20-pulse variable FA/PA train (light gray).
@@
Simulation of the fingerprint dictionary was performed in MATLAB R2020b on a Cray CS500-1211 cluster with Intel Xeon Gold 6148 high memory CPUs at the NYU Langone High Performance Computing Core facility (New York City, NY, USA). The simulation code is freely available in Matlab File Exchange (see Data Availability section for the link). Signals were simulated starting from thermal equilibrium and propagated under the optimized 23-pulse New AMRF train (Figure 1(B)). Parameter ranges ([begin:step:end]) to build the dictionary were $\text{T}_{\text{1,long}}$ = [20:2:74] ms, $\text{T}_{\text{1,short}}$ = [20:2:74] ms, $\text{T}_{\text{2,long}}^{*}$ = [10:2:66] ms, $\text{T}_{\text{2,short}}^{*}$ = [0.5:0.5:2.0, 2:2:66] ms, $\Delta\text{B}_{\text{1}}^{+}$ factor = [0.7:0.1:1.3] (applied as a multiplying factor to the RF amplitude FA) and $\Delta\text{f}_{\text{0}}$ = [-60:10:60] Hz. For $\text{T}_{\text{1}}$ quantification, we originally assumed that T1,long = T1,short. To insure that spectral densities where $J(0)\neq J(\omega)$ were included in the dictionary, a constant ($\pm\Delta\text{T}_{\text{1}}$) that added 1 ms to every entry for $\text{T}_{\text{1,long}}$ and subtracted 1 ms to every entry for $\text{T}_{\text{1,short}}$ was included in the simulations.
@@
[…]
@@
All images were reconstructed offline in MATLAB. For each channel, raw k-space data acquired during the New AMRF sequence was filtered with a Hamming kernel then reconstructed using gridding (32, 33) to produce a set of 23 complex images with a final nominal resolution $\times2.5$$\times$2.5 mm3 ($\times128$$\times$128 matrix). The images from the 8 channels were combined using coil sensitivity profiles as described by Bydder et al. (34). The average time for image reconstruction in the brain across all five volunteers was 42 s, and 38 s in the phantom. For the brain images, an additional denoising step was performed on the complex images using the Marchenko-Pastur method (35, 36, 37). This added another 15 s to the brain image reconstruction time. Supplementary Figure S1 shows 23 axial images for a center slice in the phantom. Supplementary Figures S2 and S3 show 23 reconstructed axial images for a center slice in the brain of volunteer 5.
@@ 3 Results
Relaxation time data is also presented as boxplots in Figure 4 for comparing New AMRF with RM. SD data is shown in the same figure as a comparison between New AMRF measurements and ground truth (GT), presented as mean value $\pm$5% STDV.
@@
Supplementary Figure S4 shows scatter plots with linear fits between New AMRF and RM for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$, and between New AMRF and GT (mean value $\pm$5% STDV) for SD, using data from Table 1.
@@ 4 Discussion
It is also worth noting that our New AMRF method, and the ISTO simulation that we used to generate the fingerprint dictionary, do not make any assumption about either the compartmental homogeneity within the voxel, nor the ratios between relaxation processes for the central and satellite transitions of the New Aspins $\frac{\text{3}}{\text{2}}$ (i.e. the long and short relaxatin components). Both New AMRF and the ISTO simulation simply estimate average relaxation times in each voxel, which is very likely a volume weighted average of multiple intra-voxel compartments (extracellular and intracellular spaces from multiple cell types) and local magnetic field inhomogeneities. Consequently, there is very little chance to measure the ideal ratio of 0.6:0.4 for $\text{T}_{\text{2,short}}^{*}$$\text{T}_{\text{2,long}}^{*}$ due to quadrupolar relaxation processes. Indeed, if we were to specifically measure the ratio of $\text{T}_{\text{2,short}}^{*}$$\text{T}_{\text{2,long}}^{*}$ with a fitting method (which MRF is not), the ratio of 0.6:0.4 would occur only in a perfectly ideal model where relaxation is purely quadrupolar. This is usually not the case in real samples, including gels, fluids and biological tissues. In reality, dipolar coupling, residual quadrupolar interaction (due to local structural anisotropies), local field inhomogeneities (chemical shift interaction) of various nature, and even spin-orbit interaction where the spin interacts with the magnetic fields generated by the rotational motion of the ion itself, can occur and influence the relaxation process of the New Aspins. Such influences effectively act as pertubations to the main quadrupolar relaxation process, and thus generate New Arelaxation times that deviate from the ideal case of pure quadrupolar relaxation (50, 51, 52).

### v2

@@ 2 Material and Methods
$\begin{split}\hat{\Gamma}\{\rho-\rho^{th}\}=A\sum^{2}_{m=-2}(-1)^{m}[T_{2,m},[T_{2,-m},\rho-\rho^{th}]]\\ \times\big{(}J(m\omega)-iK(m\omega)\big{)},\end{split}\vspace{-4pt}$ (2)
@@
When simulating the spin dynamics for New AMRF pulse train optimization or for generating the final fingerprint dictionary, the algorithm takes the relaxation times as input, then calculates the spectral density functions according to Equations 3-6 and uses the results to construct the Redfield relaxation superoperator in Equation 2, which is then added to the Liouville equation. After each time step of the simulation (100 $\mu$s), the simulated New AMR signal, which corresponds to the MR-observable transverse magnetization, is calculated as the average rank-1 single quantum coherence $T_{1,-1}=\frac{1}{\sqrt{2}}I_{-}=\frac{1}{\sqrt{2}}(I_{x}-iI_{y})$ using the standard formula $\langle T_{1,-1}\rangle=\text{Tr}(\rho T_{1,-1})$, where $\text{Tr}(A)$ is the trace of matrix $A$.
Figure 1: Pulse sequence diagram for 3D New AMRF. The overall pulse scheme is shown in (A). The diagram in (B) represents the variable flip angle (FA) and phase angle (PA) MRF train. There were a total of $N$ = 23 non-selective rectangular RF pulses of duration $\tau_{RF}=0.8$ ms. Each pulse was followed by a time period $\tau_{i}$. On the RF channel in (A), the gray rectangles correspond to a single FA = $\theta_{i}$ and PA = $\phi_{i}$ combination in train (B). In (B), the dark gray rectangles represent a 3-pulse inversion composite block (25) used to increase $\text{T}_{\text{1}}$ sensitivity of the sequence and improve RF homogeneity for the magnetization inversion. Within the composite block, $\tau_{1}$ = $\tau_{2}$ = 7.5 ms. The light gray rectangles indicate the 20-pulse variable FA and PA train. All FA = $\theta_{i}$ and PA = $\phi_{i}$ are listed at the top of each rectangle and $\tau_{i}$ = 15 ms for $i\geq 3$. During each delay $\tau_{i}$, a time period TE = 0.2 ms was followed by the ADC event, indicated by a yellow block on the ADC channel. A 3D spiral encoding scheme using the FLORET trajectory (20, 24, 26) (3 hubs at 45, 100 interleaves/hub), indicated in blue, was played over the gradient channels. Immediately after the ADC, a rewind gradient was also played out and then the next RF pulse in the MRF train was initiated. After $N$ pulses and $N\tau_{i}$ delays have played out, a 5-ms spoiling gradient at 70% maximum gradient strength, indicated by an orange trapezoid, was applied in all directions. The entirety of this scheme represented one TR of the sequence.
@@
Figure 1(A) presents the New AMRF sequence. System excitation was driven by a series of $N$ non-selective rectangular RF pulses with FA $\theta_{i}$ and PA $\phi_{i}$ followed by a delay $\tau_{i}$ ($i=1,2,...,N$), forming the MRF pulse train shown in Figure 1(B). An initial magnetization inversion using a 90–180–90 (dark gray) composite 3-pulse block (25) was used to increase $\text{T}_{\text{1}}$ sensitivity and improve RF homogeneity for the inversion before initiating the 20-pulse variable FA/PA train (light gray).
@@
Simulation of the fingerprint dictionary was performed in MATLAB R2020b on a Cray CS500-1211 cluster with Intel Xeon Gold 6148 high memory CPUs at the NYU Langone High Performance Computing Core facility (New York City, NY, USA). The simulation code is freely available in Matlab File Exchange (see Data Availability section for the link). Signals were simulated starting from thermal equilibrium and propagated under the optimized 23-pulse New AMRF train (Figure 1(B)). Parameter ranges ([begin:step:end]) to build the dictionary were $\text{T}_{\text{1,long}}$ = [20:2:74] ms, $\text{T}_{\text{1,short}}$ = [20:2:74] ms, $\text{T}_{\text{2,long}}^{*}$ = [10:2:66] ms, $\text{T}_{\text{2,short}}^{*}$ = [0.5:0.5:2.0, 2:2:66] ms, $\Delta\text{B}_{\text{1}}^{+}$ factor = [0.7:0.1:1.3] (applied as a multiplying factor to the RF amplitude FA) and $\Delta\text{f}_{\text{0}}$ = [-60:10:60] Hz. For $\text{T}_{\text{1}}$ quantification, we originally assumed that T1,long = T1,short. To insure that spectral densities where $J(0)\neq J(\omega)$ were included in the dictionary, a constant (±$\pm\Delta\text{T}_{\text{1}}$) that added 1 ms to every entry for $\text{T}_{\text{1,long}}$ and subtracted 1 ms to every entry for $\text{T}_{\text{1,short}}$ was included in the simulations.
@@
For reference, a 1H MPRAGE was acquired with 1 average, TR = 2300 ms, TE = 2.84 ms, FOV = 256$\times$216 mm2, slice thickness = 1 mm, 176 slices/slab and GRAPPA acceleration factor 2, for a total acquisition time of 4:32 min. The average total session time for calibration, shimming and scanning across all volunteers was 55 $\pm$ 3 min.
@@
All images were reconstructed offline in MATLAB. For each channel, raw k-space data acquired during the New AMRF sequence was filtered with a Hamming kernel then reconstructed using gridding (32, 33) to produce a set of 23 complex images with a final nominal resolution 2.5$\times$2.5$\times$2.5 mm3 (128$\times$128$\times$128 matrix). The images from the 8 channels were combined using coil sensitivity profiles as described by Bydder et al. (34). The average time for image reconstruction in the brain across all five volunteers was 42 s, and 38 s in the phantom. For the brain images, an additional denoising step was performed on the complex images using the Marchenko-Pastur method (35, 36, 37). This added another 15 s to the brain image reconstruction time. Supplementary Figure S1 shows 23 axial images for a center slice in the phantom. Supplementary Figures S2 and S3 show 23 reconstructed axial images for a center slice in the brain of volunteer 5.
@@ 3 Results
Relaxation time data is also presented as boxplots in Figure 4 for comparing New AMRF with RM. SD data is shown in the same figure as a comparison between New AMRF measurements and ground truth (GT), presented as mean value $\pm$5STDV.
@@
Supplementary Figure S4 shows scatter plots with linear fits between New AMRF and RM for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$, and between New AMRF and GT (mean value $\pm$5STDV) for SD, using data from Table 1.
@@ 4 Discussion
It is also worth noting that our New AMRF method, and the ISTO simulation that we used to generate the fingerprint dictionary, do not make any assumption about either the compartmental homogeneity within the voxel, nor the ratios between relaxation processes for the central and satellite transitions of the New Aspins $\frac{\text{3}}{\text{2}}$ (i.e. the long and short relaxation components). Both New AMRF and the ISTO simulation simply estimate average relaxation times in each voxel, which is very likely a volume weighted average of multiple intra-voxel compartments (extracellular and intracellular spaces from multiple cell types) and local magnetic field inhomogeneities. Consequently, there is very little chance to measure the ideal ratio of 0.6:0.4 for $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ due to quadrupolar relaxation processes. Indeed, if we were to specifically measure the ratio of $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ with a fitting method (which MRF is not), the ratio of 0.6:0.4 would occur only in a perfectly ideal model where relaxation is purely quadrupolar. This is usually not the case in real samples, including gels, fluids and biological tissues. In reality, dipolar coupling, residual quadrupolar interaction (due to local structural anisotropies), local field inhomogeneities (chemical shift interaction) of various nature, and even spin-orbit interaction where the spin interacts with the magnetic fields generated by the rotational motion of the ion itself, can occur and influence the relaxation process of the New Aspins. Such influences effectively act as pertubations to the main quadrupolar relaxation process, and thus generate New Arelaxation times that deviate from the ideal case of pure quadrupolar relaxation (50, 51, 52).